# algorithmic

### v1

@@ 4 Experiments and Results
In our experiments, we compare ARROW-Diff to five baseline methods that can scale to large graph generation. These methods cover the different graph generation techniques which are explained in Section 1 and Section 2. To evaluate the performance of all methods, we use five different real-world citation graph datasets, each containing a single undirected graph and all considered to be large with varying number of nodes/edges.
@@ 4.1 Experimental Setup
We use five citation graph datasets to evaluate our method: Cora-ML (McCallum et al., 2000), Cora (McCallum et al., 2000), CiteSeer (Giles et al., 1998), DBLP (Pan et al., 2016), and PubMed (Sen et al., 2008). For Cora-ML and Cora, we use the pre-processed version from Bojchevski & Günnemann (2018). Each of the five datasets contains one single, undirected, large-scale citation graph. Motivated by Bojchevski et al. (2018), we only take the largest connected component (LCC) of Cora-ML, Cora, CiteSeer, and DBLP, which all contain multiple connected components. Table 1 gives an overview of different characteristics for each graph/LCC. Similar to Bojchevski et al. (2018), we split the edges of each graph into training, validation, and test parts, and use only the training edges to train the baseline methods, as well as the OA-ARDM and GNN for ARROW-Diff.
Table 1: Dataset statistics of single, large-scale graph datasets used in this paper: Number of nodes, undirected edges, node features, and average node degree. A star () indicates that the statistics are reported for the LCC of the respective dataset.
@@
We use five different graph generation baseline methods, which are suitable for large graph generation to compare against ARROW-Diff: VGAE (Kipf & Welling, 2016b), Graphite (Grover et al., 2019), NetGAN Bojchevski et al. (2018), EDGE (Chen et al., 2023), and GraphRNN (You et al., 2018). Hence, previously mentioned methods like DiGress (Vignac et al., 2023) and GDSS (Jo et al., 2022), which are suitable for small graph generation, are excluded from our experiments. GraphRNN can scale to graphs with up to 5k nodes (You et al., 2018), so we only run it on the two smaller datasets Cora-ML (McCallum et al., 2000) and CiteSeer (Giles et al., 1998), since it is still affordable to train it on graphs of such sizes. The recent autoregressive diffusion-based graph generation method GraphARM (Kong et al., 2023) is also excluded due to the absence of a code repository. However, as demonstrated in the complexity analysis in Section 5, ARROW-Diff outperforms GraphARM in terms of runtime. To train the baseline methods, we use the recommended hyper-parameters from their papers and code. The training of NetGAN is performed using their proposed VAL-criterion (Bojchevski et al., 2018) for early stopping on the validation edges from the data split. The training of EDGE took between 2-4 days on each dataset. Hence, for downstream evaluation, we consider the models at epoch 2600 (CiteSeer), 5550 (Cora-ML), 250 (Cora), 450 (DBLP), and 250 (PubMed). Additionally, to fit into GPU memory, we decreased the batch size from 4 (training) and 64 (validation) to 2 to train the models on the Cora, DBLP, and PubMed datasets. The aforementioned baselines do not incorporate node features in their training, hence, for a fair comparison, we report the performance of ARROW-Diff both with and without the use of original node features. In our case, node features are used for training and inference of the GNN model.
ARROW-Diff Models Training
In the following, we provide details about the training of the OA-ARDM (Hoogeboom et al., 2022) and the GNN used in ARROW-Diff, which are trained independently. We train the OA-ARDM for random walk diffusion following the work of Hoogeboom et al. (2022), which is explained in Section 3. Specifically, we use a U-Net architecture similar to Ho et al. (2020) with one ResNet block and two levels for the down- and up-sampling processes. Similar to Bojchevski et al. (2018) we set the random walk length to $D=16$ and sample batches of random walks from the training split comprising edges from the original graph. Each node in the random walks, as well as each diffusion time step, is then represented using 64-dimensional learnable embeddings and passed as input to the U-Net. For the GNN component, we train a two-layer GCN (Kipf & Welling, 2016a) to predict the validity of edges based on perturbed versions of the training split of the input graph. Specifically, the GCN is trained by minimizing the binary cross-entropy loss between predicted edge probabilities and the ground truth labels which indicate whether an edge in the perturbed graph is valid or invalid. The predicted probabilities are computed by taking the sigmoid of the dot-product between pairs of node embeddings. As input to the GCN, we use either the original node features (ARROW-Diff with node features) or positional encodings (Vaswani et al., 2017) (ARROW-Diff without node features). The positional encodings are 64-dimensional for Cora-ML and CiteSeer, and 128-dimensional for Cora, DBLP, and PubMed. The GCN uses node embeddings of sizes 100 and 10 in the hidden and output layer, respectively. The full list of hyperparameters for training the OA-ARDMs and the GNN models can be found in the supplementary materials.
@@
4.2 Hyperparameter Tuning
ARROW-Diff generates a graph within several iterations $L$. In the following, we investigate the influence of the choice of $L$ on the performance of ARROW-Diff. In Figure 2 we show how the different graph evaluation metrics change with respect to $L$ on the CiteSeer (Giles et al., 1998) dataset. The increase trend on the average clustering coefficient, number of triangles, and maximum node degree can be explained by the increase in the total number of edges of the generated graph with respect to $L$ (see Figure 3). With increasing $L$, more edges are added, which in turn increase the maximum degrees of some nodes as well as the number of triangles. This also explains why e.g. the global clustering coefficient remains mostly constant, since the number of closed and open triangles rise simultaneously. We choose $L=10$ to generate graphs with ARROW-Diff across all datasets, since this value shows the closest results to the ground truth graph both in terms of number of triangles, and power-law exponent. In Figure 3, the strong drop in the number of edges for $L\in[1,5]$ is due to the fact that ARROW-Diff samples one random walk for every node in the graph at $L=1$, which results in a high number of edge proposals. With increasing $L$, these edges will either be selected or dropped from the final generated graph based on the sampling on the probabilities computed by the GNN model.
Table 2: Graph generation results of NetGAN (Bojchevski et al., 2018), VGAE (Kipf & Welling, 2016b), Graphite (Grover et al., 2019), EDGE (Chen et al., 2023), GraphRNN (You et al., 2018), and ARROW-Diff with and without node features on the single, large-scale graph datasets from Table 1. The performance is given in terms of the mean of six graph statistics across 10 generated graphs. The edge overlap represents the mean overlap with the edges of the original graph. The last column reports the graph generation time for all methods, which, for ARROW-Diff, is the time for executing Algorithm 2.
@@ 4.2 Hyperparameter Tuning
ARROW- 373 -0.112 5,912 1.81 0.191 0.001 57.3% 1.8
Diff
ARROW- 439 -0.117 5,821 1.82 0.182 0.001 55.2% 1.7
Diff (w/o features)
@@
ARROW- 536 -0.077 89,895 1.70 0.122 0.002 40.8% 13.7
Diff
ARROW- 579 -0.080 77,075 1.68 0.105 0.001 42.3% 13.4
Diff (w/o features)
@@
ARROW- 114 -0.192 795 2.24 0.109 0.004 57.8% 1.6
Diff
ARROW- 148 -0.178 996 2.15 0.125 0.003 63.0% 2.0
Diff (w/o features)
@@
ARROW- 478 -0.098 49,865 1.78 0.069 0.001 34.2% 11.2
Diff
ARROW- 675 -0.063 63,017 1.67 0.079 0.001 34.9% 9.4
Diff (w/o features)
@@
ARROW- 478 -0.082 44,120 1.90 0.039 0.001 42.7% 14.4
Diff
ARROW- 474 -0.126 41,379 1.85 0.034 0.001 41.2% 12.8
Diff (w/o features)
@@ 4.3 Results and Efficiency
In Table 2, we present the results for all baselines and for ARROW-Diff trained both with and without node features on the five datasets (Table 1). The results are reported as the mean across 10 generated graphs. To evaluate the stability of all methods, we also report the standard deviation of all metrics across the 10 generated graphs in Table 4. The results presented in Table 2 indicate that ARROW-Diff achieves notable improvement across many metrics, surpassing nearly all baselines in terms of triangle counts and power-law exponent. We observe that, EDGE performs the best in terms of maximum node degree across all datasets. This is due to its ability to steer the graph generation process towards a degree distribution similar to that of the original graph. We would like to highlight that, the number of edges in the graphs generated by VGAE (Kipf & Welling, 2016b) and Graphite (Grover et al., 2019) was extremely high. Thus, we consider only edges for which the predicted probability $p(A_{ij}=1|\bm{z}_{i},\bm{z}_{j})=\sigma(\bm{z}_{i}^{T}\bm{z}_{j})>0.95$ for nodes $i,j\in V$. Even with this constraint, the graphs generated by VGAE had over 50 M edges on the Cora, DBLP, and PubMed datasets, which led to an exhaustive metric computation. Therefore, we have omitted the results for these datasets in Table 2. The scalability of ARROW-Diff is demonstrated in the substantial decrease in graph generation time (down to more than $50\%$ compared to all baselines). This remains true even when generating very large graphs such as Cora, DBLP, and PubMed (column ’Time’ in Table 2). In the presented results, the graph generation time of VGAE and Graphite is almost zero because operating on tensors of sizes up to $000\times 3,$ 000 × 3 , 000 entries is not challenging for modern computers. The scalability of ARROW-Diff is further reflected in the training speed. Due to the small number of diffusion steps needed to train the random walk-based OA-ARDM (Hoogeboom et al., 2022), ARROW-Diff converges within 30 minutes on all datasets. In contrast, EDGE requires between 2-4 days on each dataset.
Figure 4: Visualization of the training graphs and generated graphs for Cora-ML and CiteSeer from NetGAN, Graphite, EDGE, and ARROW-Diff using the trained models from Section 4. We observe that ARROW-Diff is able to capture the basic structure of the ground-truth graph
@@ 4.4 Visualization of Generated Graphs
In Figure 4, we visualize the training split of the Cora-ML and CiteSeer graphs as well as one generated graph for each baseline method and for ARROW-Diff (using node features). We observe that ARROW-Diff is able to capture the basic structure of the ground truth graph. In Figure 5, we show the capability of ARROW-Diff in capturing community structures present in the community-20 dataset (Martinkus et al., 2022). To train ARROW-Diff for this setting, where the input comprises many small graphs, we train one model for each graph in the training set and then sample from all the trained models. As baselines, we choose those methods that are able to train on a dataset of multiple input graphs, like EDGE and GraphRNN. In Figure 6, we further show that, the generated graphs from ARROW-Diff have a heavy-tailed degree distribution, similar to real-world graphs.
Table 3: Runtime complexity $\mathcal{O(\cdot)}$ for all baseline models and ARROW-Diff. For all methods, $N$ is the number of nodes. For NetGAN, $D$ is the random walk length, $M$ is the number of sampled random walks, and $K$ is the average degree of the nodes. For VGAE and Graphite, $F$ is the size of the latent vector. For EDGE, $M$ is the number of edges in an input graph, and $K$ is the number of active nodes.
@@
NetGAN No Sequential $\mathcal{O}(MD+NK)$
@@
ARROW-Diff Yes Sequential $\mathcal{O}(L(ND+|E|))$
5 Complexity Analysis
ARROW-Diff can be used for fast generation of large graphs. In the following, we analyze the runtime complexity of ARROW-Diff, i.e., the complexity for executing Algorithm 2. Let $N$ denote the number of nodes and $|E|$ the number of edges in a graph, $D$ the random walk length which in this case is equal to the number of diffusion steps, and $L$ the number of graph generation steps (iterations) of ARROW-Diff. In each iteration $l\in[1,L]$, ARROW-Diff first samples one random walk of length $D$ for each start node $n\in V_{\text{start}}$ (line 6) to compute edge proposals. This step has a time complexity of $\mathcal{O}(ND)$ because $V_{\text{start}}\subseteq V$, and $V_{\text{start}}=V$ in the first step. Next, ARROW-Diff uses a GCN Kipf & Welling (2016a), to compute the probabilities for each edge in the so far generated graph including the set of newly proposed edges (line 8), which requires a maximum of $\mathcal{O}(|E|)$ operations. The computation of the new start nodes for the next iteration requires a maximum of $\mathcal{O}(|E|)$ operations with a complexity of $\mathcal{O}(|E|)$ for computing the node degrees of $\hat{G}$ (line 12), and $\mathcal{O}(N)$ to compute the probabilities (line 14) and sample the new start nodes (line 15). Overall, for $L$ generation steps, ARROW-Diff has a run time of $\mathcal{O}(L(ND+|E|))$. This is a great reduction in complexity compared to most existing diffusion-based graph generation approaches which have a runtime complexity of $\mathcal{O}(TN^{2})$ and are therefore not suitable for large-scale graph generation. The runtime of ARROW-Diff also outperforms that of EDGE (Chen et al., 2023), which is the only diffusion-based method in the literature that tests large graph generation at the magnitude we show here. In Table 3 we provide an overview of the runtime complexities required for the graph generation process for every baseline method. We also show that ARROW-Diff has a better runtime complexity compared to the recent autoregressive diffusion-based method GraphARM (Kong et al., 2023), which has a runtime complexity of $\mathcal{O}(N^{2})$, using overall $N(N-1)/2$ operations to predict the connecting edges of each node to all previously denoised nodes across all sampling steps.

### v2

@@ 4 Experiments and Results
In our experiments, we compare ARROW-Diff to six baseline methods for graph generation that can scale to large graphs. These methods cover the different graph generation techniques which are explained in Section 1 and Section 2. To evaluate the performance of all methods, we use five different real-world citation graph datasets, each containing a single undirected graph and all considered large graphs with varying numbers of nodes/edges. We also evaluate the performance of ARROW-Diff in generating more structured and non-scale-free graphs. To this end, we create a synthetic Stochastic Block Model (SBM) (Holland et al., 1983) graph with 360 nodes and 3,824 edges. The results on the SBM graph are shown in Appendix Section C.
@@ 4.1 Experimental Setup
To evaluate ARROW-Diff, we use five citation graph datasets: Cora-ML (McCallum et al., 2000), Cora (McCallum et al., 2000), CiteSeer (Giles et al., 1998), DBLP (Pan et al., 2016), and PubMed (Sen et al., 2008). For Cora-ML and Cora, we use the pre-processed versions from Bojchevski & Günnemann (2018). Each of the five datasets contains one single, undirected, large-scale citation graph. Motivated by Bojchevski et al. (2018), we only take the largest connected component (LCC) of Cora-ML, Cora, CiteSeer, and DBLP, which contain multiple connected components. Table 1 provides an overview of different characteristics for each graph/LCC. Similar to Bojchevski et al. (2018), we split the edges of each graph into training, validation, and test parts, and use only the training edges to train the baseline methods, as well as the OA-ARDM and GNN for ARROW-Diff.
Table 1: Dataset statistics of single, large-scale graph datasets used in this paper: Number of nodes, number of undirected edges, number of node features, and average node degree. A star () indicates that the statistics are reported for the LCC of the respective dataset.
@@
ARROW-Diff Training
In the following, we provide details about the training of the OA-ARDM (Hoogeboom et al., 2022) and the GNN used in ARROW-Diff, which are trained independently. We train the OA-ARDM for random walk diffusion following the work of Hoogeboom et al. (2022), which is explained in Section 3. Specifically, we use a U-Net (Ronneberger et al., 2015) architecture similar to Ho et al. (2020) with one ResNet block and two levels for the down- and up-sampling processes. Similar to Bojchevski et al. (2018) we set the random walk length, which equals the number of diffusion steps, to $D=16$ and sample batches of random walks from the training split comprising edges from the original graph. Each node in the random walks, as well as each diffusion time step, is then represented using 64-dimensional learnable embeddings and passed as input to the U-Net. For the GNN component, we train a two-layer GCN (Kipf & Welling, 2016a) to predict the validity of edges based on perturbed versions of the training split of the input graph. Specifically, the GCN is trained by minimizing the binary cross-entropy loss between predicted edge probabilities and the ground truth labels which indicate whether an edge in the perturbed graph is valid or invalid. The predicted probabilities are computed by taking the sigmoid of the dot-product between pairs of node embeddings. We ensure that the GCN model is well-tuned through: (1) The use of a perturbed version of the original graph for the training of the GCN, and (2) the use of a validation split and early stopping on the validation loss to ensure adequate training. As input to the GCN, we use either the original node features (ARROW-Diff with node features) or positional encodings (Vaswani et al., 2017) (ARROW-Diff without node features). The positional encodings are 64-dimensional for Cora-ML and CiteSeer, and 128-dimensional for Cora, DBLP, and PubMed. The GCN uses node embeddings of sizes 100 and 10 in the hidden and output layer, respectively. The full list of hyper-parameters for training the OA-ARDMs and the GNN models can be found in our code repository https://github.com/marsico-lab/arrow-diff.
@@
To compare against ARROW-Diff, we use six different graph generation baseline methods, which are suitable for large graph generation and can be trained on single-graph datasets: VGAE (Kipf & Welling, 2016b), Graphite (Grover et al., 2019), NetGAN Bojchevski et al. (2018), GraphRNN (You et al., 2018), BiGG (Dai et al., 2020), and EDGE (Chen et al., 2023). Hence, previously mentioned methods like DiGress (Vignac et al., 2023) and GDSS (Jo et al., 2022), which are suitable for small graph generation, are excluded from our experiments. GraphRNN can scale to graphs with up to 5k nodes (You et al., 2018), so we only run it on the two smaller datasets Cora-ML (McCallum et al., 2000) and CiteSeer (Giles et al., 1998), since it is still affordable to train it on graphs of such sizes. The recent autoregressive diffusion-based graph generation method GraphARM (Kong et al., 2023) is also excluded due to the absence of a code repository. However, as demonstrated in the complexity analysis in Section 5, ARROW-Diff outperforms GraphARM in terms of runtime. To train the baseline methods, we use the recommended hyper-parameters from their papers and code. The training of NetGAN is performed using their proposed VAL-criterion (Bojchevski et al., 2018) for early stopping on the validation edges from the data split. Some models for EDGE were not showing convergence even after 4 days of training. Hence, for evaluation, we considered the models at epochs 2600 (CiteSeer), 5550 (Cora-ML), 250 (Cora), 450 (DBLP), and 250 (PubMed). Additionally, to fit into GPU memory, we decreased the batch size from 4 to 2 to train on the Cora, DBLP, and PubMed datasets. We also decreased the embedding dimension parameter for BiGG from 256 to 128 for PubMed and to 64 for Cora and DBLP, to fit into memory of a single GPU. Some of the aforementioned baselines do not incorporate node features in their training, hence, for a fair comparison, we report the performance of ARROW-Diff both with and without the use of original node features. In our case, node features are used for training and inference of the GNN model.
@@
4.2 Hyper-parameter Tuning
ARROW-Diff generates a graph within several iterations $L$. In the following, we investigate the influence of the choice of $L$ on the performance of ARROW-Diff. In Figure 2 we show how the different graph evaluation metrics change with respect to $L$ on the CiteSeer (Giles et al., 1998) dataset. The increase trend on the average clustering coefficient, number of triangles, and maximum node degree can be explained by the increase in the total number of edges of the generated graph with respect to $L$ (see Figure 3). With increasing $L$, more edges are added, which in turn increase the maximum degrees of some nodes as well as the number of triangles. This also explains why e.g. the global clustering coefficient remains mostly constant, since the number of closed and open triangles rise simultaneously. We choose $L=10$ to generate graphs with ARROW-Diff across all datasets, since this value shows the closest results to the ground truth graph both in terms of number of triangles and power-law exponent. In Figure 3, the strong drop in the number of edges for $L\in[1,5]$ is due to the fact that ARROW-Diff samples one random walk for every node in the graph at $L=1$, which results in a high number of edge proposals. With increasing $L$, these edges will either be selected or dropped from the final generated graph based on the sampling on the probabilities computed by the GNN model.
@@ 4.2 Hyper-parameter Tuning
Table 2: Graph generation results of NetGAN (Bojchevski et al., 2018), VGAE (Kipf & Welling, 2016b), Graphite (Grover et al., 2019), EDGE (Chen et al., 2023), GraphRNN (You et al., 2018), BiGG (Dai et al., 2020), and ARROW-Diff with and without node features on the single, large-scale graph datasets from Table 1. The performance is given in terms of the mean of six graph statistics across 10 generated graphs. The edge overlap represents the mean overlap with the edges of the original graph. The last column reports the graph generation time for all methods, which, for ARROW-Diff, is the time for executing Algorithm 2.
@@
BiGG 65 0.030 280 1.98 0.026 0.008 1.0% 47.2
ARROW-Diff 373 -0.112 5,912 1.81 0.191 0.001 57.3% 1.8
ARROW-Diff 439 -0.117 5,821 1.82 0.182 0.001 55.2% 1.7
(w/o features)
@@
BiGG 222 -0.086 2,479 1.50 0.007 0.000 0.6% 820.4
ARROW-Diff 536 -0.077 89,895 1.70 0.122 0.002 40.8% 13.7
ARROW-Diff 579 -0.080 77,075 1.68 0.105 0.001 42.3% 13.4
(w/o features)
@@
BiGG 76 -0.198 284 2.44 0.096 0.005 3.2% 29.6
ARROW-Diff 114 -0.192 795 2.24 0.109 0.004 57.8% 1.6
ARROW-Diff 148 -0.178 996 2.15 0.125 0.003 63.0% 2.0
(w/o features)
@@
BiGG 71 0.183 795 1.60 0.004 0.001 1.4% 754.3
ARROW-Diff 478 -0.098 49,865 1.78 0.069 0.001 34.2% 11.2
ARROW-Diff 675 -0.063 63,017 1.67 0.079 0.001 34.9% 9.4
(w/o features)
@@
BiGG 54 -0.183 5 2.74 0.000 0.000 0.0% 265.7
ARROW-Diff 478 -0.082 44,120 1.90 0.039 0.001 42.7% 14.4
ARROW-Diff 474 -0.126 41,379 1.85 0.034 0.001 41.2% 12.8
(w/o features)
@@ 4.3 Results and Efficiency
In Table 2, we present the results for all baselines and for ARROW-Diff trained both with and without node features on the five datasets (Table 1). The results are reported as the mean across 10 generated graphs. To evaluate the stability of all methods, we also report the standard deviation of all metrics across the 10 generated graphs in Table 4. The results presented in Table 2 indicate that ARROW-Diff achieves notable improvement across many metrics, surpassing nearly all baselines in terms of triangle counts and power-law exponent (see generated node-degree distribution in Figure 7(a)). We observe that EDGE performs the best in terms of maximum node degree across all datasets. This is due to its ability to steer the graph generation process towards a degree distribution similar to that of the original graph. We would like to highlight that the number of edges in the graphs generated by VGAE (Kipf & Welling, 2016b) and Graphite (Grover et al., 2019) was extremely high. Thus, we consider only edges for which the predicted probability $p(A_{ij}=1|\bm{z}_{i},\bm{z}_{j})=\sigma(\bm{z}_{i}^{T}\bm{z}_{j})>0.95$ for nodes $i,j\in V$. Even with this constraint, the graphs generated by VGAE had over 50M edges on the Cora, DBLP, and PubMed datasets, which led to an exhaustive metric computation. Therefore, we have omitted the results for these datasets from Table 2. The scalability of ARROW-Diff is demonstrated in the substantial decrease in graph generation time (down to more than $50\%$ compared to all baselines). This remains true even when generating very large graphs such as Cora, DBLP, and PubMed (column ’Time’ in Table 2). In the presented results, the graph generation time of VGAE and Graphite is almost zero because operating on tensors of sizes up to $000\times 3,$ 000 × 3 , 000 entries is not challenging for modern computers. The scalability of ARROW-Diff is further reflected in the training speed. Due to the small number of diffusion steps needed to train the random walk-based OA-ARDM (Hoogeboom et al., 2022), ARROW-Diff converges within 30 minutes on all datasets. In contrast, EDGE requires between 2-4 days on each dataset.
@@ 4.4 Visualization of Generated Graphs
In Figure 4, we visualize the training split of the Cora-ML and CiteSeer graphs as well as one generated graph from NetGAN, Graphite, EDGE, BiGG, and ARROW-Diff (using node features). We observe that ARROW-Diff is able to capture the basic structure of the ground truth graph and seems to generate significantly fewer edges than most of the baseline methods.
Figure 4: Visualization of the training graphs and generated graphs for Cora-ML and CiteSeer from NetGAN, Graphite, EDGE, BiGG, and ARROW-Diff using the trained models from Table 2. We observe that ARROW-Diff is able to capture the basic structure of the original graph.
5 Complexity Analysis
ARROW-Diff can be used for fast generation of large graphs. In the following, we analyze the runtime complexity of ARROW-Diff, i.e., the complexity for executing Algorithm 2. Let $N$ denote the number of nodes and $|E|$ the number of edges in a graph, $D$ the random walk length which in this case is equal to the number of diffusion steps, and $L$ the number of graph generation steps (iterations) of ARROW-Diff. In each iteration $l\in[1,L]$, ARROW-Diff first samples one random walk of length $D$ for each start node $n\in V_{\text{start}}$ (line 6) to compute edge proposals. This step has a time complexity of $\mathcal{O}(ND)$ because $V_{\text{start}}\subseteq V$, and $V_{\text{start}}=V$ in the first step. Next, ARROW-Diff uses a GCN (Kipf & Welling, 2016a), to compute the probabilities for each edge in the so far generated graph including the set of newly proposed edges (line 8), which requires a maximum of $\mathcal{O}(|E|)$ operations. The computation of the new start nodes for the next iteration requires a maximum of $\mathcal{O}(|E|)$ operations with a complexity of $\mathcal{O}(|E|)$ for computing the node degrees of $\hat{G}$ (line 12), and $\mathcal{O}(N)$ to compute the probabilities (line 14) and sample the new start nodes (line 15). Overall, for $L$ generation steps, ARROW-Diff has a runtime of $\mathcal{O}(L(ND+|E|))$. This is a great reduction in complexity compared to most existing diffusion-based graph generation approaches which have a runtime complexity of $\mathcal{O}(TN^{2})$ and are therefore not suitable for large-scale graph generation. The runtime of ARROW-Diff also outperforms that of EDGE (Chen et al., 2023), which is the only diffusion-based method in the literature that performs large graph generation at the magnitude we show here. We also show that ARROW-Diff has a better runtime complexity compared to the recent autoregressive diffusion-based method GraphARM (Kong et al., 2023), which has a runtime complexity of $\mathcal{O}(N^{2})$, using overall $N(N-1)/2$ operations to predict the connecting edges of each node to all previously denoised nodes across all sampling steps. In Table 3 we compare the runtime complexity of the graph generation process for ARROW-Diff against that of each baseline method.
Table 3: Runtime complexity $\mathcal{O(\cdot)}$ for all baseline models and ARROW-Diff. For all methods, $N$ is the number of nodes and $M$ is the number of edges in an input graph. For NetGAN, $D$ is the random walk length, $R$ is the number of sampled random walks, and $K$ is the average degree of the nodes. For VGAE and Graphite, $F$ is the size of the latent vector. For EDGE, $K$ is the number of active nodes.
@@ 5 Complexity Analysis
NetGAN No Sequential $\mathcal{O}(RD+NK)$
@@
BiGG No Sequential $\mathcal{O}((N+M)\log(N))$
ARROW-Diff Yes Sequential $\mathcal{O}(L(ND+M))$
6 Limitations
While ARROW-Diff demonstrates promising results for large-scale graph generation, several limitations must be considered. First, the runtime complexity for inference of the GNN component ($\mathcal{O}(|E|)$) with increasing graph size might present a bottleneck, limiting the efficiency and scalability of ARROW-Diff to very large graphs. Another limitation of ARROW-Diff is that it generates graphs with the same number of nodes as the original graph, due to the OA-ARDM framework (Hoogeboom et al., 2022). Addressing these limitations in future work could provide a more comprehensive evaluation of ARROW-Diff’s performance and scalability across graphs with various sparsity and configurations.
Potential future work could focus on an adaptation of ARROW-Diff for learning on multiple graphs which would enable the evaluation of ARROW-Diff on a wider range of graph structures. Future experiments should also incorporate more comprehensive hyper-parameter tuning for the reported baselines. Relying on default hyper-parameters for each method may potentially undermine their performance. Finally, investigating the effect of the random walk length on the results could provide valuable insights into ARROW-Diff’s sensitivity to this parameter.